OpenStudy (anonymous):

Solve the proportion. 1. 3/x = 4/ (x-8) 2. m/(m-8) = 5/3

OpenStudy (shoba):

we need to cross multiply... 1) 3/x = 4/ (x-8) 3(x-8) = 4x 3x-24 = 4x x =- 24 2) m/(m-8) = 5/3 3m = 5m - 40 5m-2m = 40 2m = 40 m = 20